# Quantum Dynamic Programming

### v1

@@ IV Analysis of the nested fixed-point Grover search
Figure 1: Illustration comparing exact and QDP implementation of nested fixed-point Grover search recursions. The top two rows describe the exact recursion. The uppermost row shows $N$ recursion steps by the unitary channel $\hat{\mathbf{\Gamma}}^{(\psi_{n-1})}_{L}$ containing $L$ memory-calls to the instruction state $\psi_{n-1}$. The resulting state of $n$th recursion step is $\psi_{n}$. The second row from the top gives a more detailed picture of each recursion step, where reflections around the fixed state $\hat{\mathbf{E}}^{(\tau)}_{\beta_{i}}$ are interlaced with memory-calls $\hat{\mathbf{E}}^{(\psi_{n-1})}_{\alpha_{i}}$. The bottom three rows represent the QDP implementation. The middle one among them outlines the overall protocol of $N$ recursions, where interferential mixedness reduction (IMR) protocols follow each dynamic implementation $\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}$ approximating $\hat{\mathbf{\Gamma}}^{(\rho_{n-1})}_{L}$. Note that the instruction state $\rho_{n-1}$ for each recursion is different from the exact instruction $\psi_{n-1}$ due to the inevitable error occurring from QDP implementation. In the lowermost row, it is shown that each memory-call $\hat{\mathbf{E}}^{(\psi_{n-1})}_{\alpha_{i}}$ is implemented by $\frac{M}{L}$ DME memory-usage queries consuming $\frac{M}{L}$ copies of the instruction state. The resulting state of the approximate channel $\hat{\mathmybb{\Gamma}}^{(\rho_{n-1})}_{L}$ is $\rho_{n}^{\prime}$, which is different from the next instructions state $\rho_{n}$. The next instruction state is prepared from $\rho_{n}^{\prime}$ using IMR protocols, described in the third row from the top. Each round of IMR protocol halves the number of copies one have, and total $R$ rounds are executed at each recursion step.
In this section, we delineate the entire protocol for the QDP implementation of nested fixed-point Grover search recursions. See Fig. 1 for the overall structure of the section. The exact recursion (two uppermost rows in the figure) is analyzed in Sec. IV.1 with the explicit formula giving the distance between the resulting state and the target state after each recursion. QDP implementation of the recursion is carried out by combining two subroutines: approximation of the recursion unitary using DME memory-usage queries (lowermost row in the figure) and the interferential mixedness reduction (IMR) subroutine (third row from the top) that maintains the mixedness of the instruction states below some threshold. Both of these subroutines are introduced in Sec. IV.2. In the rest of the section, we calculate the total cost of this implementation, namely the depth of the circuit and the number of copies needed, and show that the final state can be made arbitrarily close to the desired final state.
@@ IV.1 Exact implementation of nested fixed-point Grover search
$T_{m}(x)=\begin{cases}\cos\bigl{(}m\arccos(x)\bigr{)},&\lvert x\rvert\leq 1\ ,\\ \cosh\bigl{(}m\operatorname{arcosh}(x)\bigr{)},&\lvert x\rvert>1\ .\end{cases}$ (54)
@@
$\hat{\Gamma}^{(\psi_{n-1})}_{L}=\hat{E}^{(\psi_{n-1})}_{\alpha_{L}}\hat{E}^{(\tau)}_{\beta_{L}}\cdots\hat{E}^{(\psi_{n-1})}_{\alpha_{1}}\hat{E}^{(\tau)}_{\beta_{1}}$ (56)
to the iterated state $|{\psi_{n-1}}\rangle$ obtaining $|{\psi_{n}}\rangle=\hat{\Gamma}^{(\psi_{n-1})}_{L}|{\psi_{n-1}}\rangle$. The composite unitary operation in Eq. (56) consists of partial reflection unitaries defined as
$\hat{E}^{(\psi)}_{s}=\mathmybb{1}-(1-e^{-is})|{\psi}\rangle\!\langle{\psi}|\ ,$ (57)
with angle $s$, where we have adapted the notation $\psi=|{\psi}\rangle\!\langle{\psi}|$ and $\tau=|{\tau}\rangle\!\langle{\tau}|$ for pure states. When $s=\pi$, we obtain the usual reflection unitary around the state $|{\psi}\rangle$. In Eq. (56), following in the steps of  [11], we consider $L$ partial reflections around both $|{\tau}\rangle$ and $|{\psi_{n-1}}\rangle$ with angles
$\alpha_{l}(q)=-\beta_{L-l+1}(q)=2\cot^{-1}\left(\tan\left(\frac{2\pi l}{2L+1}\right)\sqrt{1-\left[T_{1/(2L+1)}(q)\right]^{-2}}\right)\ ,$ (58)
@@
$\frac{1}{2}\left\lVert\tau-\hat{\Gamma}^{(\psi)}_{L}(\psi)\left(\hat{\Gamma}^{(\psi)}_{L}\right)^{\dagger}\right\rVert_{1}=qT_{2L+1}\left(T_{1/(2L+1)}(q^{-1})\frac{1}{2}\lVert\tau-\psi\rVert_{1}\right)\ .$ (59)
@@
Suppose that the parameter $q$ in Eq. (58) is chosen to be $q=\delta_{n}$ for the recursion unitary $\hat{\Gamma}^{(\psi_{n-1})}_{L}$ at any $n$, where
$\delta_{n}=\operatorname{sech}\left((2L+1)\operatorname{arcsech}\left(\delta_{n-1}\right)\right)\ ,\quad\forall n>0\ ,$ (60)
starting from the base $\delta_{0}=\frac{1}{2}\lVert\tau-\psi_{0}\rVert_{1}$. Then $\frac{1}{2}\left\lVert\tau-\psi_{n}\right\rVert_{1}=\delta_{n}$ for all $n$, where $\psi_{n}$ is recursively defined as $|{\psi_{n}}\rangle=\hat{\Gamma}^{(\psi_{n-1})}_{L}|{\psi_{n-1}}\rangle$.
@@
We first observe that Eq. (60) is equivalent to
$\delta_{n}^{-1}=T_{2L+1}\left(\delta_{n-1}^{-1}\right)\ ,$ (61)
@@
$T_{1/(2L+1)}(\delta_{n}^{-1})=T_{1/(2L+1)}\left(T_{2L+1}\left(\delta_{n-1}^{-1}\right)\right)=\delta_{n-1}^{-1}.$ (62)
We may now use this identity in simplifying the trace distance of interest: starting from the general relation Eq. (59) with $q=\delta_{n}$,
$\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}=\delta_{n}T_{2L+1}\left(T_{1/(2L+1)}\bigl{(}\delta_{n}^{-1}\bigr{)}\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}\right)=\delta_{n}T_{2L+1}\left(\delta_{n-1}^{-1}\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}\right)\ ,$ (63)
where the second equality comes from applying Eqs. (62). If the Lemma is true for step $n-1$, i.e. if $\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}=\delta_{n-1}$, then Eq. (63) leads to $\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}=\delta_{n}$ using $T_{2L+1}(1)=1$. Since the base case $\frac{1}{2}\lVert\tau-\psi_{0}\rVert_{1}=\delta_{0}$ is true, the lemma holds by induction. ∎
The explicit form of the $\delta_{N}$, i.e. the distance to the target state after $N$ iterations can also be derived. From Eqs. (55) and (61), $\delta_{N}^{-1}=T_{(2L+1)^{N}}\left(\delta_{0}^{-1}\right)$, or equivalently,
$\delta_{N}=\operatorname{sech}\left((2L+1)^{N}\operatorname{arcosh}\Bigl{(}\delta_{0}^{-1}\Bigr{)}\right)\ .$ (64)
For large $N$, Eq. (64) asymptotically behaves as $\delta_{N}=\mathcal{O}\left(e^{-\operatorname{arcosh}\bigl{(}\delta_{0}^{-1}\bigr{)}(2L+1)^{N}}\right)$. A typical setting of a Grover search starts from the initial fidelity $F_{0}=1-\delta_{0}^{2}\ll 1$, e.g. the initial state is a superposition of all basis states. Then the decay rate can be rewritten as
$\operatorname{arcosh}\left(\delta_{0}^{-1}\right)=\operatorname{arcosh}((1-F_{0})^{-1/2})=\sqrt{F_{0}}+\mathcal{O}(F_{0}^{3/2})\ ,$ (65)
@@
$(2L+1)^{N}=\mathcal{O}\left(\log\left(\delta_{N}^{-1}\right)F_{0}^{-1/2}\right)$ (66)
is needed for the final distance $\delta_{N}$. The circuit depth for the entire algorithm is proportional to Eq. (66) due to the unfolding implementation, and it contains the characteristic Grover scaling factor $F_{0}^{-1/2}$. If a non-recursive algorithm is applied, Eq. (64) can still be used with $N=1$. Denote the final distance as
$\delta^{\prime}=\operatorname{sech}\left((2L^{\prime}+1)\operatorname{arcosh}\left(\delta_{0}^{-1}\right)\right),$ (67)
when the alternation length of this algorithm is $L^{\prime}$. For $\delta^{\prime}=\delta_{N}$, the alternation length of the non-nested algorithm should satisfy $2L^{\prime}+1=(2L+1)^{N}$, resulting in exponentially many partial reflections. Note that Eq. (66) then implies $L^{\prime}\sim F_{0}^{-1/2}$, which displays the quantum advantage compared to the classical search requiring $L^{\prime}\sim F_{0}^{-1}$.
@@ IV.2.1 Distance bounds after a DME implementation
In this subsection, we introduce the DME implementation of an iteration of the dynamic Grover search. We first establish (through Proposition 11 and Corollary 12) that the DME implementation still has the essential feature of the Grover search, namely all relevant states are effectively two-dimensional. Then the main results Lemma 14 and Corollary 15 provide the upper bound of a distance after one iteration of DME implemented Grover search, starting from a mixed state. In short, we observe that whenever the initially prepared state is close to being pure, then the distance to our Grover target state still follows a similar recursive relation as before, namely Eq. (91).
The ‘dynamic’ parts of the recursion unitary $\hat{\Gamma}^{(\rho)}_{L}$ are the memory-calls $\hat{E}^{(\rho)}_{\alpha}$, i.e. partial reflectors around the state $\rho$. Hence, DME (Def. 6) is the right choice of memory-usage query to use for implementing the memory-call $\hat{E}^{(\rho)}_{\alpha}$. To recap, the DME memory-usage query can be written as
^s$\hat{\mathcal{E}}^{(\rho)}_{s}\left(\sigma\right)=\cos^{2}(s)\sigma-i\sin(s)\cos(s)[\rho,\sigma]+\sin^{2}(s)\rho\ ,$ (68)
with the duration $s$ of our choice. If $M$ queries are made with $s=\frac{\alpha}{M}$, the implemented channel (^s$(\hat{\mathcal{E}}^{(\rho)}_{s}(\sigma))^{M}$ approximates the memory-call $\hat{E}^{(\rho)}_{\alpha}$ with error $\mathcal{O}(1/M)$.
@@
    The implementation of a recursion step $\hat{\Gamma}^{(\psi)}_{L}$ becomes non-unitary even when the instruction state $\psi=|{\psi}\rangle\!\langle{\psi}|$ and the input state $\phi=|{\phi}\rangle\!\langle{\phi}|$ are pure. To see this, consider a special case of DME protocol with pure instruction state $\psi$ and input state $\phi$,
    ^$\hat{\mathcal{E}}^{(\psi)}_{s}(\phi)=\cos^{2}(s)\phi+\sin^{2}(s)\psi-i\sin(s)\cos(s)\langle{\psi}|{\phi}\rangle|{\psi}\rangle\!\langle{\phi}|+\mathrm{h.c.}.$ (69)
@@
    $\operatorname{Tr}\left[\left\{\hat{\mathcal{E}}^{(\psi)}_{s}(\phi)\right\}^{2}\right]$ $=\cos^{4}(s)+\sin^{4}(s)+2\sin^{2}(s)\cos^{2}(s)\left(2\lvert\langle{\psi}|{\phi}\rangle\rvert^{2}-\lvert\langle{\psi}|{\phi}\rangle\rvert^{4}\right)$ (70)
@@
    $\mathcal{H}_{\rm rel}\coloneq\mathrm{span}\left\{|{\tau}\rangle,|{\tau^{\perp}}\rangle\coloneq\frac{|{\psi_{0}}\rangle-\langle{\tau}|{\psi_{0}}\rangle|{\tau}\rangle}{\sqrt{1-|\langle{\tau}|{\psi_{0}}\rangle|^{2}}}\right\},$ (71)
@@
    $\mathcal{S}_{\rm rel}\coloneq\left\{\rho\,|\,\rho:\text{density matrix},\ \mathrm{supp}(\rho)\subset\mathcal{H}_{\rm rel}\right\}$ (72)
@@
Suppose two states $\rho,\sigma\in\mathcal{S}_{\rm rel}$. The state after a DME channel $\sigma^{\prime}=\hat{\mathcal{E}}^{(\rho)}_{s}(\sigma)$ defined in Eq. (39) is also contained in $\mathcal{S}_{\rm rel}$. In other words, for all $s\in\mathbb{R}$, we have that ^s$\hat{\mathcal{E}}^{(\mathcal{S}_{\mathrm{rel}})}_{s}(\mathcal{S}_{\mathrm{rel}})\subset\mathcal{S}_{\mathrm{rel}}$.
@@
^s$\hat{\mathcal{E}}^{(\rho)}_{s}(\sigma)|{\phi}\rangle=\cos^{2}(s)\sigma|{\phi}\rangle-i\sin(s)\cos(s)[\rho,\sigma]|{\phi}\rangle+\sin^{2}(s)\rho|{\phi}\rangle=0\ ,$ (73)
whenever $|{\phi}\rangle\in\mathcal{H}_{\rm rel}^{\mathsf{C}}$. Therefore, $\mathrm{supp}(\hat{\mathcal{E}}^{(\rho)}_{s}(\sigma))\subset\mathcal{H}_{\rm rel}$. ∎
@@
From Proposition 11, DME protocols map states in $\mathcal{S}_{\rm rel}$ to $\mathcal{S}_{\rm rel}$. Furthermore, reflections around the target state $\tau\in\mathcal{S}_{\rm rel}$, which is implemented by an exact unitary, not DME, also ensures that $\hat{\mathbf{E}}^{(\tau)}_{s}\left(\sigma\right)\in\mathcal{S}_{\rm rel}$ whenever $\sigma\in\mathcal{S}_{\rm rel}$. Since a recursion step of Grover search comprises the above two operations, the state after the iteration, $\rho_{n}$, also stays in $\mathcal{S}_{\rm rel}$. ∎
@@
$\hat{\mathbf{\Gamma}}^{(\phi)}_{L}\left(\mathmybb{1}_{\rm rel}\right)=\mathmybb{1}_{\rm rel},$ (74)
due to unitarity of $\hat{\mathbf{\Gamma}}^{(\phi)}_{L}$. Since $\dim(\mathcal{H}_{\rm rel})=2$, we can decompose any $\rho\in\mathcal{S}_{\rm rel}$ in its eigenbasis
$\rho=(1-a)|{\psi}\rangle\!\langle{\psi}|+a|{\psi^{\perp}}\rangle\!\langle{\psi^{\perp}}|=(1-x)|{\psi}\rangle\!\langle{\psi}|+x\frac{\mathmybb{1}_{\rm rel}}{2},$ (75)
for $x=2a\in[0,1]$ by using the fact that $\mathmybb{1}_{\rm rel}=|{\psi}\rangle\!\langle{\psi}|+|{\psi^{\perp}}\rangle\!\langle{\psi^{\perp}}|$. We will refer to $\psi$ as the pure state associated with $\rho$ and $x$ as the mixedness parameter. Interestingly, the partial reflection around $\rho$ with angle $s$ can be rewritten using the observation that $[\psi,\mathmybb{1}_{\rm rel}]=0$:
$\hat{E}^{(\rho)}_{s}=e^{-is\left[(1-x)\psi+x\frac{\mathmybb{1}_{\rm rel}}{2}\right]}=e^{-isx\frac{\mathmybb{1}_{\rm rel}}{2}}\hat{E}^{(\psi)}_{s(1-x)}\ ,$ (76)
which is identical to the partial reflection around the corresponding pure state $\psi$ with the angle reduced by the factor $(1-x)$, apart from the phase factor globally acting on $\mathcal{H}_{\rm rel}$. If the mixedness parameter $x$ is known, one can effectively implement the reflection around the pure state $\psi$ using mixed states $\rho$ by adjusting the DME duration $s\mapsto s/(1-x)$.
Nevertheless, it is usually the case that only an upper bound, but not the exact amount of mixedness $x$ is known. The Proposition and the Lemma below assess the QDP implementation of a Grover search recursion step when the mixedness parameter $x$ is unknown. Instead of the exact partial reflection unitary channels $\hat{\mathbf{E}}^{(\psi)}_{s}=\hat{\mathbf{E}}^{(\rho)}_{s/(1-x)}$ around its associated pure state $\psi$, we employ DME queries ^s$\hat{\mathcal{E}}^{(\rho)}_{s}$ as defined in Eq. (68).

### v2

@@ IV Analysis of the nested fixed-point Grover search
In this section, we delineate the entire protocol for the QDP implementation of nested fixed-point Grover search recursions, using it as an illustrative example. The exact recursion is analyzed in Sec. IV.1 with the explicit formula giving the distance between the resulting state and the target state after each recursion. QDP implementation of the recursion is carried out by combining two subroutines: approximation of the recursion unitary using DME memory-usage queries and the interferential mixedness reduction (IMR) subroutine that maintains the mixedness of the instruction states below some threshold. The first subroutine is analysed in detail in Sec. IV.2. In the rest of the section, we calculate the total cost of this implementation, namely the depth of the circuit and the number of copies needed, and show that the final state can be made arbitrarily close to the desired final state.
@@ IV.1 Exact implementation of nested fixed-point Grover search
$T_{m}(x)=\begin{cases}\cos\bigl{(}m\arccos(x)\bigr{)},&\lvert x\rvert\leq 1\ ,\\ \cosh\bigl{(}m\operatorname{arcosh}(x)\bigr{)},&\lvert x\rvert>1\ .\end{cases}$ (S70)
@@
$\hat{\Gamma}^{[\psi_{n-1}]}_{L}=\hat{E}^{[\psi_{n-1}]}_{\alpha_{L}}\hat{E}^{[\tau]}_{\beta_{L}}\cdots\hat{E}^{[\psi_{n-1}]}_{\alpha_{1}}\hat{E}^{[\tau]}_{\beta_{1}}$ (S72)
to the iterated state $|{\psi_{n-1}}\rangle$ obtaining $|{\psi_{n}}\rangle=\hat{\Gamma}^{[\psi_{n-1}]}_{L}|{\psi_{n-1}}\rangle$. The composite unitary operation in Eq. (S72) consists of partial reflection unitaries defined as
$\hat{E}^{[\psi]}_{s}\coloneq\mathmybb{1}-(1-e^{-is})|{\psi}\rangle\!\langle{\psi}|\ ,$ (S73)
with angle $s$, where we have adapted the notation $\psi=|{\psi}\rangle\!\langle{\psi}|$ and $\tau=|{\tau}\rangle\!\langle{\tau}|$ for pure states. When $s=\pi$, we obtain the usual reflection unitary around the state $|{\psi}\rangle$. In Eq. (S72), following in the steps of  [5], we consider $L$ partial reflections around both $|{\tau}\rangle$ and $|{\psi_{n-1}}\rangle$ with angles
$\alpha_{l}(q)=-\beta_{L-l+1}(q)=2\cot^{-1}\left(\tan\left(\frac{2\pi l}{2L+1}\right)\sqrt{1-\left[T_{1/(2L+1)}(q)\right]^{-2}}\right)\ ,$ (S74)
@@
$\frac{1}{2}\left\lVert\tau-\hat{\Gamma}^{[\psi]}_{L}(\psi)\left(\hat{\Gamma}^{[\psi]}_{L}\right)^{\dagger}\right\rVert_{1}=qT_{2L+1}\left(T_{1/(2L+1)}(q^{-1})\frac{1}{2}\lVert\tau-\psi\rVert_{1}\right)\ .$ (S75)
@@
Suppose that the parameter $q$ in Eq. (S74) is chosen to be $q=\delta_{n}$ for the recursion unitary $\hat{\Gamma}^{[\psi_{n-1}]}_{L}$ at any $n$, where
$\delta_{n}=\operatorname{sech}\left((2L+1)\operatorname{arcsech}\left(\delta_{n-1}\right)\right)\ ,\quad\forall n>0\ ,$ (S76)
starting from the base $\delta_{0}=\frac{1}{2}\lVert\tau-\psi_{0}\rVert_{1}$. Then $\frac{1}{2}\left\lVert\tau-\psi_{n}\right\rVert_{1}=\delta_{n}$ for all $n$, where $\psi_{n}$ is recursively defined as $|{\psi_{n}}\rangle=\hat{\Gamma}^{[\psi_{n-1}]}_{L}|{\psi_{n-1}}\rangle$.
@@
We first observe that Eq. (S76) is equivalent to
$\delta_{n}^{-1}=T_{2L+1}\left(\delta_{n-1}^{-1}\right)\ ,$ (S77)
@@
$T_{1/(2L+1)}(\delta_{n}^{-1})=T_{1/(2L+1)}\left(T_{2L+1}\left(\delta_{n-1}^{-1}\right)\right)=\delta_{n-1}^{-1}.$ (S78)
We may now use this identity in simplifying the trace distance of interest: starting from the general relation Eq. (S75) with $q=\delta_{n}$,
$\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}=\delta_{n}T_{2L+1}\left(T_{1/(2L+1)}\bigl{(}\delta_{n}^{-1}\bigr{)}\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}\right)=\delta_{n}T_{2L+1}\left(\delta_{n-1}^{-1}\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}\right)\ ,$ (S79)
where the second equality comes from applying Eqs. (S78). If the Lemma is true for step $n-1$, i.e. if $\frac{1}{2}\lVert\tau-\psi_{n-1}\rVert_{1}=\delta_{n-1}$, then Eq. (S79) leads to $\frac{1}{2}\lVert\tau-\psi_{n}\rVert_{1}=\delta_{n}$ using $T_{2L+1}(1)=1$. Since the base case $\frac{1}{2}\lVert\tau-\psi_{0}\rVert_{1}=\delta_{0}$ is true, the lemma holds by induction. ∎
The explicit form of the $\delta_{N}$, i.e. the distance to the target state after $N$ iterations can also be derived. From Eqs. (S71) and (S77), $\delta_{N}^{-1}=T_{(2L+1)^{N}}\left(\delta_{0}^{-1}\right)$, or equivalently,
$\delta_{N}=\operatorname{sech}\left((2L+1)^{N}\operatorname{arcosh}\Bigl{(}\delta_{0}^{-1}\Bigr{)}\right)\ .$ (S80)
For large $N$, Eq. (S80) asymptotically behaves as $\delta_{N}=\mathcal{O}\left(e^{-\operatorname{arcosh}\bigl{(}\delta_{0}^{-1}\bigr{)}(2L+1)^{N}}\right)$. A typical setting of a Grover search starts from the initial fidelity $F_{0}=1-\delta_{0}^{2}\ll 1$, e.g. the initial state is a superposition of all basis states. Then the decay rate can be rewritten as
$\operatorname{arcosh}\left(\delta_{0}^{-1}\right)=\operatorname{arcosh}((1-F_{0})^{-1/2})=\sqrt{F_{0}}+\mathcal{O}(F_{0}^{3/2})\ ,$ (S81)
@@
$(2L+1)^{N}=\mathcal{O}\left(\log\left(\delta_{N}^{-1}\right)F_{0}^{-1/2}\right)$ (S82)
is needed for the final distance $\delta_{N}$. The circuit depth for the entire algorithm is proportional to Eq. (S82) due to the unfolding implementation, and it contains the characteristic Grover scaling factor $F_{0}^{-1/2}$. If a non-recursive algorithm is applied, Eq. (S80) can still be used with $N=1$. Denote the final distance as
$\delta^{\prime}=\operatorname{sech}\left((2L^{\prime}+1)\operatorname{arcosh}\left(\delta_{0}^{-1}\right)\right),$ (S83)
when the alternation length of this algorithm is $L^{\prime}$. For $\delta^{\prime}=\delta_{N}$, the alternation length of the non-nested algorithm should satisfy $2L^{\prime}+1=(2L+1)^{N}$, resulting in exponentially many partial reflections. Note that Eq. (S82) then implies $L^{\prime}\sim F_{0}^{-1/2}$, which displays the quantum advantage compared to the classical search requiring $L^{\prime}\sim F_{0}^{-1}$.
@@ IV.2.1 Distance bounds after a DME implementation
In this subsection, we introduce the DME implementation of an iteration of the dynamic Grover search. We first establish (through Proposition 14 and Corollary 15) that the DME implementation still has the essential feature of the Grover search, namely all relevant states are effectively two-dimensional. Then the main results Lemma 17 and Corollary 18 provide the upper bound of a distance after one iteration of DME implemented Grover search, starting from a mixed state. In short, we observe that whenever the initially prepared state is close to being pure, then the distance to our Grover target state still follows a similar recursive relation as before, namely Eq. (S99).
The ‘dynamic’ parts of the recursion unitary $\hat{\Gamma}^{[\rho]}_{L}$ are the memory-calls $\hat{E}^{[\rho]}_{\alpha}$, i.e. partial reflectors around the state $\rho$. Hence, DME (Def. 6) is the right choice of memory-usage query to use for implementing the memory-call $\hat{E}^{[\rho]}_{\alpha}$. To recap, the DME memory-usage query can be written as
^s$\hat{\mathcal{E}}^{[\rho]}_{s}\left(\sigma\right)=\cos^{2}(s)\sigma-i\sin(s)\cos(s)[\rho,\sigma]+\sin^{2}(s)\rho\ ,$ (S84)
with the duration $s$ of our choice. If $M$ queries are made with $s=\frac{\alpha}{M}$, the implemented channel (^s$(\hat{\mathcal{E}}^{[\rho]}_{s}(\sigma))^{M}$ approximates the memory-call $\hat{E}^{[\rho]}_{\alpha}$ with error $\mathcal{O}(1/M)$.
@@
    The implementation of a recursion step $\hat{\Gamma}^{[\psi]}_{L}$ becomes non-unitary even when the instruction state $\psi=|{\psi}\rangle\!\langle{\psi}|$ and the input state $\phi=|{\phi}\rangle\!\langle{\phi}|$ are pure. To see this, consider a special case of DME protocol with pure instruction state $\psi$ and input state $\phi$,
    ^$\hat{\mathcal{E}}^{[\psi]}_{s}(\phi)=\cos^{2}(s)\phi+\sin^{2}(s)\psi-i\sin(s)\cos(s)\langle{\psi}|{\phi}\rangle|{\psi}\rangle\!\langle{\phi}|+\mathrm{h.c.}.$ (S85)
@@
    $\operatorname{Tr}\left[\left\{\hat{\mathcal{E}}^{[\psi]}_{s}(\phi)\right\}^{2}\right]$ $=\cos^{4}(s)+\sin^{4}(s)+2\sin^{2}(s)\cos^{2}(s)\left(2\lvert\langle{\psi}|{\phi}\rangle\rvert^{2}-\lvert\langle{\psi}|{\phi}\rangle\rvert^{4}\right)$ (S86)
@@
    $\mathcal{H}_{\rm rel}\coloneq\mathrm{span}\left\{|{\tau}\rangle,|{\tau^{\perp}}\rangle\coloneq\frac{|{\psi_{0}}\rangle-\langle{\tau}|{\psi_{0}}\rangle|{\tau}\rangle}{\sqrt{1-|\langle{\tau}|{\psi_{0}}\rangle|^{2}}}\right\},$ (S87)
@@
    $\mathcal{S}_{\rm rel}\coloneq\left\{\rho\,|\,\rho:\text{density matrix},\ \mathrm{supp}(\rho)\subset\mathcal{H}_{\rm rel}\right\}$ (S88)
@@
Suppose two states $\rho,\sigma\in\mathcal{S}_{\rm rel}$. The state after a DME channel $\sigma^{\prime}=\hat{\mathcal{E}}^{[\rho]}_{s}(\sigma)$ defined in Eq. (S26) is also contained in $\mathcal{S}_{\rm rel}$. In other words, for all $s\in\mathbb{R}$, we have that ^s$\hat{\mathcal{E}}^{[\mathcal{S}_{\mathrm{rel}}]}_{s}(\mathcal{S}_{\mathrm{rel}})\subset\mathcal{S}_{\mathrm{rel}}$.
@@
^s$\hat{\mathcal{E}}^{[\rho]}_{s}(\sigma)|{\phi}\rangle=\cos^{2}(s)\sigma|{\phi}\rangle-i\sin(s)\cos(s)[\rho,\sigma]|{\phi}\rangle+\sin^{2}(s)\rho|{\phi}\rangle=0\ ,$ (S89)
whenever $|{\phi}\rangle\in\mathcal{H}_{\rm rel}^{\mathsf{C}}$. Therefore, $\mathrm{supp}(\hat{\mathcal{E}}^{[\rho]}_{s}(\sigma))\subset\mathcal{H}_{\rm rel}$. ∎
@@
From Proposition 14, DME protocols map states in $\mathcal{S}_{\rm rel}$ to $\mathcal{S}_{\rm rel}$. Furthermore, reflections around the target state $\tau\in\mathcal{S}_{\rm rel}$, which is implemented by an exact unitary, not DME, also ensures that $\hat{\mathbf{E}}^{[\tau]}_{s}\left(\sigma\right)\in\mathcal{S}_{\rm rel}$ whenever $\sigma\in\mathcal{S}_{\rm rel}$. Since a recursion step of Grover search comprises the above two operations, the state after the iteration, $\rho_{n}$, also stays in $\mathcal{S}_{\rm rel}$. ∎
@@
$\hat{\mathbf{\Gamma}}^{[\phi]}_{L}\left(\mathmybb{1}_{\rm rel}\right)=\mathmybb{1}_{\rm rel},$ (S90)
due to unitarity of $\hat{\mathbf{\Gamma}}^{[\phi]}_{L}$. Since $\dim(\mathcal{H}_{\rm rel})=2$, we can decompose any $\rho\in\mathcal{S}_{\rm rel}$ in its eigenbasis
$\rho=(1-x)|{\psi}\rangle\!\langle{\psi}|+x|{\psi^{\perp}}\rangle\!\langle{\psi^{\perp}}|=(1-2x)|{\psi}\rangle\!\langle{\psi}|+x\mathmybb{1}_{\rm rel},$ (S91)
for $x\in[0,\frac{1}{2}]$ by using the fact that $\mathmybb{1}_{\rm rel}=|{\psi}\rangle\!\langle{\psi}|+|{\psi^{\perp}}\rangle\!\langle{\psi^{\perp}}|$. We will refer to $\psi$ as the pure state associated with $\rho$ and $x$ as the mixedness parameter. Interestingly, the partial reflection around $\rho$ with angle $s$ can be rewritten using the observation that $[\psi,\mathmybb{1}_{\rm rel}]=0$:
$\hat{E}^{[\rho]}_{s}=e^{-is\left[(1-2x)\psi+x\mathmybb{1}_{\rm rel}\right]}=e^{-isx\mathmybb{1}_{\rm rel}}\hat{E}^{[\psi]}_{s(1-2x)}\ ,$ (S92)
which is identical to the partial reflection around the corresponding pure state $\psi$ with the angle reduced by the factor $(1-2x)$, apart from the phase factor globally acting on $\mathcal{H}_{\rm rel}$. If the mixedness parameter $x$ is known, one can effectively implement the reflection around the pure state $\psi$ using mixed states $\rho$ by adjusting the DME duration $s\mapsto s/(1-2x)$.
Nevertheless, it is usually the case that only an upper bound, but not the exact amount of mixedness $x$ is known. The Proposition and the Lemma below assess the QDP implementation of a Grover search recursion step when the mixedness parameter $x$ is unknown. Instead of the exact partial reflection unitary channels $\hat{\mathbf{E}}^{[\psi]}_{s}=\hat{\mathbf{E}}^{[\rho]}_{s/(1-2x)}$ around its associated pure state $\psi$, we employ DME queries ^s$\hat{\mathcal{E}}^{[\rho]}_{s}$ as defined in Eq. (S84).